OpenStudy (anonymous):

For which system of equations is (–2, 2) the solution? A. 3x – 5y = –16 –2x + 3y = 10 B. –4x + 2y = 6 2x – 3y = 10 C. 4x + 3y = 2 5x + 4y = –2 D. 6x + 5y = –2 –8x + 5y = –6

OpenStudy (solomonzelman):

plug in the solution into your equation. (-2,2) ( compare to (x,y) ) plug in -2 for x, and 2 for y into each system of equations.

OpenStudy (anonymous):

@Tennistar @amistre64

OpenStudy (the_fizicx99):

You could solve each individually or just plug in (-2,2) for x and y and see which gives you a true statement.

OpenStudy (anonymous):

oh duh its c

OpenStudy (the_fizicx99):

o.o how did you get c

OpenStudy (the_fizicx99):

4x + 3y = 2 4(-2) + 3(2) = 2 -8 + 6 = -2 -2 = 2 <-- not true

OpenStudy (anonymous):

oh so thats why i messed up? so its B @tHe_FiZiCx99

OpenStudy (the_fizicx99):

Show me how you got B?